# Coarse geometry of quasi-transitive graphs beyond planarity

### v3

@@ 1. Introduction
    for each $e=uv$ there is a path $P_{e}$ whose two endpoints lie in $M_{u}$ and $M_{v}$ and internal vertices are not in $\bigcup_{v\in V(H)}M_{v}$, and
@@
A graph is locally finite if every vertex has finite degree. Georgakopoulos and Papasoglu conjectured the following [GP23].
@@
The main technical tool that we use is a recent structural theorem on locally finite quasi-transitive graphs excluding the countable clique as a minor [EGLD23], which shows that such graphs have a canonical tree-decomposition in which all torsos are planar or finite (see the next section for the definitions). Most importantly, this result does not use the Robertson-Seymour graph minor structure theorem.
@@
We can view each graph $G$ as a metric space, by considering the natural shortest-path metric associated to $G$.
It was proved in [${}^{+}$20] that every bounded degree graph excluding a minor has asymptotic dimension at most 2, and that any planar graph has asymptotic dimension at most 2. This was improved in [${}^{+}$23], where it was shown that any graph excluding a minor has asymptotic dimension at most 2, and any planar graph has Assouad-Nagata dimension at most 2. This was finally extended by Liu in [Liu23], who proved that any graph avoiding a minor has Assouad-Nagata dimension at most 2 (a different proof was then given by Distel in [Dis23]). All the results on graphs excluding a minor mentioned above (even for bounded degree graphs) crucially rely on the Graph minor structure theorem by Robertson and Seymour [RS03], a deep result proved in a series of 16 papers.
Using the invariance of Assouad-Nagata dimension under bilipschitz embedding [LS05], we will give a short proof of the fact that minor-excluded quasi-transitive graphs have Assouad-Nagata dimension at most 2. A finitely generated group is said to be minor-excluded if it has a Cayley graph which excludes a minor. Our result directly implies that finitely generated minor-excluded groups have Assouad-Nagata dimension at most 2, and thus asymptotic dimension at most 2 (which was originally conjectured by Ostrovskii and Rosenthal in [OR15]). We will only use Theorem 1.3 and a few simple tools from [${}^{+}$20, ${}^{+}$23] based on the work of Brodskiy, Dydak, Levin and Mitra [BDLM08]. In particular we will give a short proof of the fact that planar graphs of bounded degree have Assouad-Nagata dimension at most 2. Crucially, our proof for quasi-transitive graphs excluding a minor does not rely on the Graph minor structure theorem of Robertson and Seymour.
@@ 2. Preliminaries
The sets $V_{t}\cap V_{t^{\prime}}$ for every $tt^{\prime}\in E(T)$ are called the adhesion sets of $(T,\mathcal{V})$ and the adhesion of $(T,\mathcal{V})$ is the maximum size of its adhesion sets (possibly infinite). For $t\in V(T)$, the torso $G\llbracket V_{t}\rrbracket$ is the graph obtained from $G[V_{t}]$ (the subgraph of $G$ induced by the bag $V_{t}$) by adding all edges $uv$ with $u,v\in V_{t}$ for which there exist $t^{\prime}$ such that $u$ and $v$ lie in the adhesion set $V_{t}\cap V_{t^{\prime}}$.
@@
We will need the main result of [EGLD23], which gives the structure of locally finite quasi-transitive graph excluding a minor (note that this result does not use the Graph minor structure theorem by Robertson and Seymour [RS03]).
Theorem 2.1 ([EGLD23]).
@@ 4. Beyond planarity
We let $f:V(G)\to V(H)$ and $A>0$ be such that for each $x,x^{\prime}\in V(G)$:
@@
and such that the $A$-neighborhood of $f(V(G))$ covers $H$. We also let $\Delta\in\mathbb{N}$ denote the maximum degree of $G$. Note that for each $xy\in E(G)$, there exists a $f(x)f(y)$-path $P_{xy}$ in $H$ such that $|P_{xy}|\leqslant A^{2}$. We let $H^{\prime}$ be the subgraph of $H$ given by the union of all such paths $P_{xy}$.
We first observe that $H^{\prime}$ is quasi-isometric to $G$, and that $f$ gives the corresponding quasi-isometric embedding. Note that for every $z\in V(H^{\prime})$, by construction there must be some edge $xy\in E(G)$ such that $z\in P_{xy}$. In particular, $d_{H^{\prime}}(z,f(x))\leqslant A^{2}$. Note that by construction we clearly have $d_{H^{\prime}}(f(x),f(y))\leqslant A^{2}d_{G}(x,y)$ for each $x,y\in V(G)$, and as $d_{H^{\prime}}(f(x),f(y))\geqslant d_{H}(x,y)$, $f$ induces indeed a quasi-isometric embedding between $G$ and $H^{\prime}$.
Now we show that $H^{\prime}$ has bounded degree. Let $z\in V(H^{\prime})$ and $xy\in E(G)$ such that $z\in V(P_{xy})$. Then $d_{H^{\prime}}(z,f(x))\leqslant A^{2}$ so $X:=\{x\in V(G),z\in V(P_{xy})\}$ has diameter at most $2A^{2}$ in $G$. Note that as $G$ degree at most $\Delta$, we have $|X|\leqslant\Delta^{2A^{2}}$. In particular it implies that $H^{\prime}$ has degree at most $\Delta^{2A^{2}}$. ∎
Given a graph $G$ and an integer $k\geqslant 1$, the $k$-th power of $G$, denoted by $G^{k}$, is the graph with the same vertex set as $G$ in which two vertices are adjacent if and only if they are at distance at most $k$ in $G$. The $k$-blow up of $G$, denoted by $G\boxtimes K_{k}$, is the graph obtained from $G$ by replacing each vertex $u$ by a copy $C_{u}$ of the complete graph $K_{k}$, and by adding all edges between pairs $C_{u},C_{v}$ if and only if $u$ and $v$ are adjacent in $G$ (so that each edge of $G$ is replaced by a complete bipartite graph $K_{k,k}$ in $G\boxtimes K_{k}$). Quasi-isometries of bounded degree graphs are related to graph powers and blow-ups by the following lemma.
@@
We let $A>0$ and $f:V(G)\to V(H)$ be such that for each $x,x^{\prime}\in V(G)$:
@@
By Lemmas 4.2 and 4.3, there is a graph $H_{1}$ of bounded local crossing number and maximum degree and an integer $\ell$ such that $G$ is a subgraph of $H_{1}^{\ell}\boxtimes K_{\ell}$. By Observation 4.4, there is graph $H_{2}$ of bounded local crossing number and maximum degree such that $G$ is a subgraph of $H_{2}^{\ell+2}$. Observe that every $s$-planar graph $F_{1}$ is a subgraph of $F_{2}^{s+1}$, where $F_{2}$ is the planar graph obtained from $F_{1}$ by placing a new vertex at each crossing (and note that if $F_{1}$ has bounded degree, then $F_{2}$ also has bounded degree). It follows that there is a planar graph $H$ of bounded degree and an integer $k$, such that $G$ is a subgraph of $H^{k}$. ∎
@@ 5. Assouad-Nagata dimension of minor-excluded groups
When such a function $f$ exists we also say that the graph $G$ has a bilipschitz embedding in $H$. Since graphs are uniformly discrete metric spaces (in the sense that any two distinct elements lie at distance at least 1 apart), any quasi-isometric embedding of a graph $G$ in a graph $H$ is also a bilipschitz embedding of $G$ in $H$. Hence, Theorem 1.3 has the following immediate consequence.
@@
We first prove that planar graphs of bounded degree have Assouad-Nagata dimension at most 2 (a stronger version of this result, without the bounded degree assumption, was proved in [${}^{+}$23]).
We need the following result, first proved in [BST12] in a slightly different form. Another proof can be found in [${}^{+}$20, ${}^{+}$23] based on a result of Ding and Oporowski [DO95] which states that every graph of treewidth at most $t$ and maximum degree at most $\Delta$ is a subgraph of the strong product of a tree with a complete graph on $24t\Delta$ vertices. The proofs of all these results are fairly short.
Theorem 5.2 ([BST12]).
If a graph $G$ has bounded degree and bounded treewidth, then $G$ has Assouad-Nagata dimension at most 1.
A layering of a graph $G$ is a partition of the vertex set of $G$ into sets $L_{0},L_{1},\ldots$, called layers, so that any pair of adjacent vertices in $G$ either lies in the same layer or in consecutive layers (i.e. layers $L_{i},L_{i+1}$ for some $i\geqslant 1$). A simple example of layering is given by a BFS-layering of $G$, obtained by choosing one root vertex $v_{C}$ in each connected component $C$ of $G$, and then defining $L_{i}$ (for all $i\geqslant 0$) as the set of vertices of $G$ at distance exactly $i$ from one of the vertices $v_{C}$.
It was proved by Bodlaender [Bod88] that planar graphs of bounded diameter have bounded treewidth. This directly implies the following.
Lemma 5.3.
For any BFS-layering of a planar graph $G$, and any integer $k$, the subgraph of $G$ induced by $k$ consecutive layers of $\mathcal{L}$ has bounded treewidth.
Proof.
Let $L_{0},L_{1},\ldots$ be a BFS-layering of $G$. Consider the planar subgraph $H$ of $G$ induced by $k$ consecutive layers $L_{i},L_{i+1},\ldots,L_{i+k-1}$. If $i=0$, then $H$ is a disjoint union of graphs of radius at most $k$, and thus has bounded treewidth by Bodlaender’s result [Bod88]. Assume now that $i\geqslant 1$, and let $H^{+}$ be the supergraph of $H$ obtained by adding a vertex $r$ that dominates all the vertices of $L_{i}$. Note that $H^{+}$ can be obtained from the subgraph of $G$ induced by the layers $L_{0},L_{1},\ldots,L_{i+k-1}$, by contracting all layers $L_{0},L_{1},\ldots,L_{i-1}$ (which induce a connected subgraph of $G$, by definition of a BFS-layering) into a single vertex. Thus $H^{+}$ is a minor of $G$, which implies that $H^{+}$ is planar. Moreover, it follows from the definition of a BFS-layering that each vertex of $H^{+}$ lies at distance at most $k$ from $r$, and thus $H^{+}$ has diameter at most $2k$. By Bodlaender’s result [Bod88], $H^{+}$ has bounded treewidth, and thus $H$ (as a subgraph of $H^{+}$) also has bounded treewidth. ∎
The next result appears as Theorem 4.3 in [${}^{+}$23], and is a simple application of the main result in [BDLM08] (which has a nice and short combinatorial proof).
Theorem 5.4 ([${}^{+}$23]).
If a graph $G$ has a layering $\mathcal{L}=(L_{0},L_{1},\ldots)$ such that for any integer $k$, the disjoint union of all subgraphs of $G$ induced by $k$ consecutive layers of $\mathcal{L}$ has Assouad-Nagata dimension at most $n$, then $G$ has Assouad-Nagata dimension at most $n+1$.
We immediately deduce the following.
@@
Every planar graph $G$ of bounded degree has Assouad-Nagata dimension at most 2.
Proof.
Consider a BFS-layering $\mathcal{L}$ of $G$. By Lemma 5.3, for any $k\geqslant 1$, the disjoint union of all subgraphs of $G$ induced by $k$ consecutive layers of $\mathcal{L}$ has bounded treewidth. As $G$ has bounded degree, this disjoint union of subgraphs of $G$ also has bounded degree, and thus by Theorem 5.2 it has Assouad-Nagata dimension at most 1. Hence, it follows from Theorem 5.4 that $G$ itself has Assouad-Nagata dimension at most 2. ∎
We are now ready to prove the main result of this section. Recall that a stronger version (without the quasi-transitivity assumption) was proved in [Liu23, Dis23], but the version below has a reasonably simple proof that does not rely on the Robertson-Seymour graph minor structure theorem.
Theorem 5.6.
@@
Proof.
By Corollary 5.1, every locally finite quasi-transitive $K_{\infty}$-minor free graph $G$ has a bilipschitz embedding in some planar graph $H$ of bounded degree, which has Assouad-Nagata dimension at most 2. Since Assouad-Nagata dimension is invariant under bilipschitz embedding [LS05], $G$ has Assouad-Nagata dimension at most 2. ∎
@@ 6. Open problems
It was proved by MacManus [Mac23] that if a finitely generated group has a Cayley graph which is quasi-isometric to a planar graph, then it is quasi-isometric to a planar Cayley graph. In the same spirit, it is natural to ask whether we can require the planar graph of bounded degree in Theorem 1.3 to be quasi-transitive, or even a Cayley graph. Our current proof does preserve symmetries, as we do a number of non-canonical choices for the images of the vertices.
@@
Here is perhaps a more interesting example. The strong product $G\boxtimes H$ of two graphs $G$ and $H$ has vertex set $V(G)\times V(H)$, and two distinct vertices $(u,x)$ and $(v,y)$ are adjacent if and only if ($u=v$ or $uv\in E(G)$) and ($x=y$ or $xy\in E(H)$). Consider the strong product $T\boxtimes P$ of the infinite binary tree $T$ and the infinite 2-way path $P$. Using Theorems 5.2 and Lemma 5.4, this graph has Assouad-Nagata dimension at most 2. As it contains a quasi-isometric copy of a 2-dimensional grid, the Assouad-Nagata dimension of $T\boxtimes P$ is indeed equal to 2. On the other hand, we observe that for any integer $k$, $T\boxtimes P$ contains the complete bipartite graph $K_{k,k}$ as an asymptotic minor. To see this, remark that $T$ contains an infinite $k$-claw (the graph obtained by gluing $k$ infinite 1-way paths at their starting vertex) as an asymptotic minor (obtained by contracting a subtree of $T$ with $k$ leaves into a single vertex, and pruning the additional branches). The strong product of this infinite $k$-claw with $P$ consists of $k$ copies of an infinite 2-dimensional grid (restricted to the upper half-plane, say), glued on a common infinite path $\pi$. On this path we can select $k$ vertices, arbitrarily far apart, and on each infinite grid we can select a single vertex, arbitrarily far from $\pi$, and connect it to the $k$ vertices of $\pi$ via disjoint paths. By taking a ball of sufficiently large radius around each of the $2k$ vertices, we obtain $K_{k,k}$ as an $r$-fat minor for arbitrarily large $r$, and thus $K_{k,k}$ as an asymptotic minor. This is illustrated for $k=3$ in Figure 1. Since containing a graph $H$ as an asymptotic minor is invariant under quasi-isometry, any graph $G$ which is quasi-isometric to $T\boxtimes P$ also contains every $K_{k,k}$ as an asymptotic minor, and thus every finite graph as a minor (in particular, $G$ cannot be planar).

### v6

@@ 1. Introduction
    for each $e=uv\in E(H)$ there is a path $P_{e}$ whose two endpoints lie in $M_{u}$ and $M_{v}$ and internal vertices are not in $\bigcup_{v\in V(H)}M_{v}$, and
@@
We can view each graph $G$ as a metric space, by considering the natural shortest-path metric associated to $G$. A graph is locally finite if every vertex has finite degree. Georgakopoulos and Papasoglu conjectured the following [GP23].
@@
The main technical tool that we use is a recent structural theorem on locally finite quasi-transitive graphs excluding the countable clique as a minor [EGLD24], which shows that such graphs have a canonical tree-decomposition in which all torsos are planar or finite (see the next section for the definitions). Most importantly, this result does not use the Robertson-Seymour graph minor structure theorem.
@@
It was first proved that planar graphs have Assouad-Nagata dimension at most 3 [FP21], later improved to 2 by [BBE+24, JL22]. It was also proved in [BBE+24] that any graph excluding a minor has asymptotic dimension at most 2. This was finally extended by Liu in [Liu23], who proved that any graph avoiding a minor has Assouad-Nagata dimension at most 2 (a different proof was then given by Distel in [Dis23]). All the results on graphs excluding a minor mentioned above (even earlier results in [BBE+20] for bounded degree graphs) crucially rely on the Graph minor structure theorem by Robertson and Seymour [RS03], a deep result proved in a series of 16 papers.
Using the invariance of Assouad-Nagata dimension under bilipschitz embedding [LS05], we will deduce from Theorem 1.3 that minor-excluded quasi-transitive graphs have Assouad-Nagata dimension at most 2. A finitely generated group is said to be minor-excluded if it has a Cayley graph which excludes a minor. Our result directly implies that finitely generated minor-excluded groups have Assouad-Nagata dimension at most 2, and thus asymptotic dimension at most 2 (which was originally conjectured by Ostrovskii and Rosenthal in [OR15]). Crucially, our proof does not rely on the Graph minor structure theorem of Robertson and Seymour.
@@ 2. Preliminaries
The sets $V_{t}\cap V_{t^{\prime}}$ for every $tt^{\prime}\in E(T)$ are called the adhesion sets of $(T,\mathcal{V})$ and the adhesion of $(T,\mathcal{V})$ is the supremum of the sizes of its adhesion sets (possibly infinite). For $t\in V(T)$, the torso $G\llbracket V_{t}\rrbracket$ is the graph obtained from $G[V_{t}]$ (the subgraph of $G$ induced by the bag $V_{t}$) by adding all edges $uv$ with $u,v\in V_{t}$ for which there exist $t^{\prime}$ such that $u$ and $v$ lie in the adhesion set $V_{t}\cap V_{t^{\prime}}$.
@@
We will need the main result of [EGLD24], which gives the structure of locally finite quasi-transitive graph excluding a minor (note that this result does not use the Graph minor structure theorem by Robertson and Seymour [RS03]).
Theorem 2.1 ([EGLD24]).
@@ 4. Beyond planarity
We let $f:V(G)\to V(H)$ and $A\geqslant 1$ be such that for each $x,x^{\prime}\in V(G)$:
@@
and such that the $A$-neighborhood of $f(V(G))$ covers $H$. We also let $\Delta\in\mathbb{N}$ denote the maximum degree of $G$. Note that for each $xy\in E(G)$, there exists a $f(x)f(y)$-path $P_{xy}$ in $H$ such that $|P_{xy}|\leqslant A\cdot 1+1=2A$. We let $H^{\prime}$ be the subgraph of $H$ given by the union of all such paths $P_{xy}$.
We first observe that $H^{\prime}$ is quasi-isometric to $G$, and that $f$ gives the corresponding quasi-isometric embedding. Note that for every $z\in V(H^{\prime})$, by construction there must be some edge $xy\in E(G)$ such that $z\in P_{xy}$. In particular, $d_{H^{\prime}}(z,f(x))\leqslant 2A$. Note that by construction we clearly have $d_{H^{\prime}}(f(x),f(y))\leqslant 2Ad_{G}(x,y)$ for each $x,y\in V(G)$, and as $d_{H^{\prime}}(f(x),f(y))\geqslant d_{H}(f(x),f(y))$, $f$ indeed induces a quasi-isometric embedding between $G$ and $H^{\prime}$.
Now we show that $H^{\prime}$ has bounded degree. Let $z\in V(H^{\prime})$ and $xy\in E(G)$ such that $z\in V(P_{xy})$. Then $d_{H^{\prime}}(z,f(x))\leqslant 2A$ so $X:=\{x\in V(G),z\in V(P_{xy})\}$ has diameter at most $4A$ in $G$. Note that as $G$ degree at most $\Delta$, we have $|X|\leqslant\Delta^{4A}$. In particular it implies that $H^{\prime}$ has degree at most $\Delta^{4A}$. ∎
Given a graph $G$ and an integer $k\geqslant 1$, the $k$-th power of $G$, denoted by $G^{k}$, is the graph with the same vertex set as $G$ in which two vertices are adjacent if and only if they are at distance at most $k$ in $G$. The $k$-blow-up of $G$, denoted by $G\boxtimes K_{k}$, is the graph obtained from $G$ by replacing each vertex $u$ by a copy $C_{u}$ of the complete graph $K_{k}$, and by adding all edges between pairs $C_{u},C_{v}$ if and only if $u$ and $v$ are adjacent in $G$ (so that each edge of $G$ is replaced by a complete bipartite graph $K_{k,k}$ in $G\boxtimes K_{k}$). Quasi-isometries of bounded degree graphs are related to graph powers and blow-ups by the following lemma.
@@
We let $A\geqslant 1$ and $f:V(G)\to V(H)$ be such that for each $x,x^{\prime}\in V(G)$:
@@
By Lemmas 4.2 and 4.3, there is a graph $H_{1}$ of bounded local crossing number and maximum degree and an integer $\ell$ such that $G$ is a subgraph of $H_{1}^{\ell}\boxtimes K_{\ell}$. By Observation 4.4, there is a graph $H_{2}$ of bounded local crossing number and maximum degree such that $G$ is a subgraph of $H_{2}^{\ell+2}$. Observe that every $s$-planar graph $F_{1}$ is a subgraph of $F_{2}^{s+1}$, where $F_{2}$ is the planar graph obtained from $F_{1}$ by placing a new vertex at each crossing (and note that if $F_{1}$ has bounded degree, then $F_{2}$ also has bounded degree). It follows that there is a planar graph $H$ of bounded degree and an integer $k$, such that $G$ is a subgraph of $H^{k}$. ∎
@@ 5. Assouad-Nagata dimension of minor-excluded groups
When such a function $f$ exists we also say that the graph $G$ has a bilipschitz embedding in $H$. Since graphs are uniformly discrete metric spaces (in the sense that any two distinct elements lie at distance at least 1 apart), for any quasi-isometric embedding of a bounded degree graph $G$ in a graph $H$, there is a bilipschitz embedding of $G$ in a graph obtained from $H$ by adding, for each vertex $v$ of $H$, a bounded number of vertices of degree 1 adjacent to $v$ (note that in the journal version of the paper [EG24], it is incorrectly stated that with the same assumptions there is a bilipschitz embedding of $G$ in $H$). Hence, Theorem 1.3 has the following immediate consequence.
@@
It was proved in [BBE+24, JL22] that planar graphs have Assouad-Nagata dimension at most 2. The proofs are similar and both quite short: they combine an original argument of Fujiwara and Papasoglu [FP21], proving that planar graphs have Assouad-Nagata dimension at most 3, with a dimension reduction technique introduced by Brodskiy, Dydak, Levin and Mitra [BDLM08]. Since Assouad-Nagata dimension is invariant under bilipschitz embedding [LS05], we obtain the following as a direct application.
@@
Recall that a stronger version (without the quasi-transitivity assumption) was proved in [Liu23, Dis23], but the version above has a reasonably simple proof that does not rely on the Robertson-Seymour graph minor structure theorem.
Important note
In the journal version of the paper [EG24] we gave a simple proof that planar graphs of bounded degree have Assouad-Nagata dimension at most 2. However our argument was incorrect, as we were incorrectly using [BBE+24, Theorem 4.3]. More precisely, we stated in [EG24, Theorem 5.4] that whenever any graph of a class $\mathcal{C}$ has a layering such that any number of consecutive layers induces a class of Assouad-Nagata dimension at most $n$, then the class $\mathcal{C}$ has Assouad-Nagata dimension at most $n+1$. This is false, as graphs of bounded layered treewidth show (these graphs have unbounded Assouad-Nagata dimension [BBE+24], while any bounded number of layers induce a graph of bounded treewidth, and such graphs have Assouad-Nagata dimension at most 1 [BBE+24]). Indeed, to apply [BBE+24, Theorem 4.3], it is required that in addition, the $n$-dimensional control function is also linear in the number of layers (which is not the case for graphs of bounded layered treewidth).
We would like to apologize to everyone for the inconvenience, and express our gratitude to Robert Hickingbotham, who noticed the issue.
@@ 6. Open problems
It was proved by MacManus [Mac23] that if a finitely generated group has a Cayley graph which is quasi-isometric to a planar graph, then it is quasi-isometric to a planar Cayley graph. In the same spirit, it is natural to ask whether we can require the planar graph of bounded degree in Theorem 1.3 to be quasi-transitive, or even a Cayley graph. Our current proof does not preserve symmetries, as we do a number of non-canonical choices for the images of the vertices. As remarked by a referee, the stronger question above, whether we can require the planar graph in Theorem 1.3 to be a Cayley graph, is a special case of a Problem of Woess [Woe91, Problem 1], which asked whether every transitive graph is quasi-isometric to a Cayley graph. This turned out to have a negative answer [EFW12] in general, but the question restricted to (quasi-)transitive graphs excluding a minor might still have a positive answer.
@@
Here is perhaps a more interesting example. The strong product $G\boxtimes H$ of two graphs $G$ and $H$ has vertex set $V(G)\times V(H)$, and two distinct vertices $(u,x)$ and $(v,y)$ are adjacent if and only if ($u=v$ or $uv\in E(G)$) and ($x=y$ or $xy\in E(H)$). Consider the strong product $T\boxtimes P$ of the infinite binary tree $T$ and the infinite 2-way path $P$. This graph has Assouad-Nagata dimension 2. On the other hand, we observe that for any integer $k$, $T\boxtimes P$ contains the complete bipartite graph $K_{k,k}$ as an asymptotic minor. To see this, remark that $T$ contains an infinite $k$-claw (the graph obtained by gluing $k$ infinite 1-way paths at their starting vertex) as an asymptotic minor (obtained by contracting a subtree of $T$ with $k$ leaves into a single vertex, and pruning the additional branches). The strong product of this infinite $k$-claw with $P$ consists of $k$ copies of an infinite 2-dimensional grid (restricted to the upper half-plane, say), glued on a common infinite path $\pi$. On this path we can select $k$ vertices, arbitrarily far apart, and on each infinite grid we can select a single vertex, arbitrarily far from $\pi$, and connect it to the $k$ vertices of $\pi$ via disjoint paths. By taking a ball of sufficiently large radius around each of the $2k$ vertices, we obtain $K_{k,k}$ as an $r$-fat minor for arbitrarily large $r$, and thus $K_{k,k}$ as an asymptotic minor. This is illustrated for $k=3$ in Figure 1. Since containing a graph $H$ as an asymptotic minor is invariant under quasi-isometry, any graph $G$ which is quasi-isometric to $T\boxtimes P$ also contains every $K_{k,k}$ as an asymptotic minor, and thus every finite graph as a minor (in particular, $G$ cannot be planar).